# The Giroux Correspondence in dimension 3

### v1

@@ Abstract.
In [LV] the authors proved the Giroux Correspondence for tight contact $3$-manifolds via convex Heegaard surfaces. Simultaneously, [HBH] gave an all-dimensions proof of the Giroux Correspondence by generalising convex surface theory to higher dimensions. This paper uses a key result from [HBH] and also [Tia] to complete the $3$-dimensional proof for arbitrary (not necessarily tight) contact 3-manifolds. This presentation features low-dimensional techniques and further clarifies the relationship between contact manifolds and their Heegaard splittings that was introduced in [LV].
@@ 1. Introduction
[Giroux Correspondence [Gir02]] Two open book decompositions support isotopic contact structures if and only if they admit a common positive stabilisation.
Because topology and contact geometry are so closely entwined, it’s natural to probe whether topological techniques can be squeezed to produce meaningful contact geometric outputs. In this spirit, in 2023 the authors developed new tools to study $3$-dimensional contact manifolds via their Heegaard decompositions.
We used the enhanced relationship between contact structures and Heegaard splittings to prove the Giroux Correspondence in the case of tight contact $3$-manifolds [LV]. Simultaneously, Breen-Honda-Huang used generalised convex surface theory to prove the Giroux Correspondence for all contact structures and in all dimensions [HBH]. While the work of Breen-Honda-Huang fully covers the $3$-dimensional case, the generality of their techniques lends their proof a different flavour. This paper completes the 3-dimensional story, using an essential result of [HBH] and also [Tia] to bridge the gap to overtwisted manifolds that limited the scope of our previous paper. We believe this work will make the proof more accessible to a low-dimensional audience; in particular, it offers a useful illustration of convex surface techniques in concert with Heegaard splittings.
@@ 1.1. Heegaard splittings and open book decompositions
Any open book decomposition of a manifold gives rise to a canonical Heegaard decomposition. Heegaard splittings produced thus are called convex Heegaard splittings, and one may directly recover an open book supporting $\xi$ from any convex Heegaard splitting of $(M,\xi)$. The innovation in [LV] is a procedure called refinement that produces an open book from a more general class of Heegaard splittings of a contact manifold. Although this open book decomposition is not uniquely defined, any two open books produced thus admit a common positive stabilisation – exactly the equivalence appearing in the Giroux Correspondence. It follows that the Giroux Correspondence can be rephrased as a statement about convex Heegaard splittings (Theorem 3.3). Our proof of the Giroux Correspondence for tight contact manifolds relied on showing that many operations one might naturally perform on a Heegaard splitting of a contact manifold in fact preserve the positive stabilisation class of the associated open book.
However, these techniques were insufficient in the case of overtwisted manifolds. In order to assign an open book decomposition to a contact manifold with a Heegaard splitting, our approach required both handlebodies to be tight. When two convex Heegaard surfaces are isotopic, there is a sequence of convex Heegaard surfaces interpolating between them. However, even when both original Heegaard splittings have tight handlebodies, intermediate Heegaard splittings in the sequence may nevertheless have overtwisted handlebodies. The present work bypasses this stumbling block by considering highly stabilised Heegaard splittings where the handlebodies are always tight. We introduce the term “bridging” for a particular method of constructing such a splitting, and we show that different choices made in constructing a bridge splitting preserve the positive stabilisation class of the associated open book. This allows us to prove that any pair of convex Heegaard splittings for a fixed contact manifold correspond to open books that admit a common positive stabilisation.
@@ 1.2. A note on notation
The arguments in this paper involve sequences of Heegaard splittings of varying types, each of which is denoted by some sort of “$\mathcal{H}$”. We have attempted to define notation locally and maintain these choices consistently, but we briefly outline the system here for reference. Every Heegaard surface $\Sigma$ is required to be convex; Heegaard splittings subject to no further restrictions are denoted by $\mathcal{H}=(\Sigma,U,V)$, where $U$ and $V$ are the handlebodies, oriented so that $\partial U=\Sigma$. Elements of a pair of Heegaard splittings are distinguished by the addition of a prime: $\mathcal{H}=(\Sigma,U,V)$ and $\mathcal{H}^{\prime}=(\Sigma^{\prime},U^{\prime},V^{\prime})$. Following the notation in [LV], the refinement of $\mathcal{H}$ is denoted by ~$\widetilde{\mathcal{H}}=(\widetilde{\Sigma},\widetilde{U},\widetilde{V})$.
Starting in Section 3.2, we will stabilise Heegaard splittings by attaching certain contact $1$-handles found in the complement of fixed handlebodies. We will call this process bridging, and we denote the bridge of $\mathcal{H}$ by ^$\widehat{\mathcal{H}}=(\widehat{\Sigma},\widehat{U},\widehat{V})$. When we bridge only a single handlebody of the original splitting, we denote the result by ^$\widehat{\mathcal{H}}^{U}$ or ^$\widehat{\mathcal{H}}^{V}$, as appropriate. Distinct bridged splittings may be distinguished by primes: $\widehat{\mathcal{H}}$ and $\widehat{\mathcal{H}}^{\prime}$.
@@ Acknowledgment
This research was supported in part by the Austrian Science Fund (FWF) P 34318. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. The authors would also like to thank Ko Honda and Matthias Scharitzer for helpful discussions on the topic.
@@ 2. Background
We assume the reader is familiar with basic notions in contact geometry, but we use this section to collect some tools and terminology that will be essential for the rest of the paper. Section 2.1 provides a quick review of convex surface theory and introduces weak contactomorphism (Definition 2.1) as an equivalence relation on contact manifolds with convex boundary. Section 2.2 describes the relationship between open book decompositions and convex Heegaard splittings (Definition 2.4) of a contact manifold. Section 2.3 presents the contact handle model for bypass attachment. Finally, Section 2.4 states the key theorem borrowed from [HBH] and [Tia] which classifies distinct bypass decompositions of a product cobordism. Also, Lemma 2.8 gives a new perspective on the familiar operation of bypass rotation.
@@ 2.1. Convex surfaces
The reader looking for a more thorough introduction to convex surface theory is referred to [Mas14] or the lecture notes [Etn04, Hon].
@@
[LV, Definition 2.3.] Two contact manifolds $(M,\xi)$ and $(M^{\prime},\xi^{\prime})$ with convex boundary are weakly contactomorphic if there is a contact embedding $\iota\colon M\hookrightarrow M^{\prime}$ such that $\iota(\partial M)$ is convexly isotopic to $\partial M^{\prime}$.
@@
We note that if $\iota:M\rightarrow M^{\prime}$ is a contact embeddeding as above, then $M\setminus\iota({M^{\prime}})$ is a weak identity morphism in the sense of [Tia].
Henceforth, we will be interested in contact manifolds with boundary only up to weak contactomorphism. Two (weak contactomorphism classes of) manifolds may be glued along diffeomorphic boundary components if and only if there is a diffeomorphism respecting the dividing sets. Gluing is well defined as an operation on weak contactomorphism classes of manifolds; see [LV, Proposition 2.8.].
@@
[LV, Definition 2.4.] Two embedded codimension-00 submanifolds $N_{1},N_{2}\subset(M,\xi)$ with convex boundary are weakly (contact) isotopic if there is an isotopy $N_{s}$ between them such that $\partial N_{s}$ remains convex throughout.
@@ A Heegaard splitting $(\Sigma,U,V)$ of $(M,\xi)$ is convex if
[LV, Proposition 3.4.] Let $(M,\xi)$ be a contact manifold and let $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ be open books supporting $\xi$. Then the induced convex Heegaard splittings are isotopic via an isotopy keeping the Heegaard surface convex if and only if $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ are isotopic through a path of open books supporting $\xi$.
@@ 2.3. Bypass attachment
Figure 1. Left: a bypass half-disc attached along the arc $c$. Centre: the attached $1$-handle is a regular neighbourhood of an arc Legendrian isotopic to $c$. Right: attaching the shaded $2$-handle along $c\cup c^{\prime}$ completes the bypass attachment.
[…]
@@ 2.4. Bypass decompositions
[…]
The main theorem of both [Tia] and [HBH] establishes that Trival Insertion and Far Commutation suffice to relate any two bypass decompositions of topologically trivial products:
@@
[HBH, Theorem 3.1.2 ] Any two bypass decompositions of $(\Sigma\times I,\xi,\Gamma_{\Sigma})$ are related by isotopy of the bypass arcs and a finite iteration of Trivial Insertion and Far Commutation moves.
The dimension 3 case was the Main Theorem in [Tia], although presented in different language.
@@
Figure 4. Attach a $1$-handle $h_{T}^{1}$ along a push-off of the Legendrian curve $d$. Then isotope the disc cobounded by $h_{T}^{1}$ and $d$ across the product disc $A$ to produce a bypass disc $D$ cobounded by $h_{T}^{1}$ and $c_{T}$.
@@ 3. Heegaard Splittings and the Giroux Correspondence
With useful tools and vocabulary established, we now turn to proving the Giroux Correspondence. As a first step, Section 3.1 summarises results from [LV] that allow us to restate the theorem about a pair of open books as a theorem about a pair of Heegaard splittings. Section 3.2 introduces bridging as an operation to stabilise a Heegaard splitting of a contact manifold. After establishing some properties of bridge splittings, we relate them to refinements in Section 3.3 to prove the main result.
@@ 3.1. Tight Heegaard splittings and refinement
As described in Section 2.2, each convex Heegaard splitting of a contact manifold corresponds to an open book decomposition supporting the contact structure. Here we recall a key idea from [LV] that extends the class of Heegaard splittings of a contact manifold which determine open books.
@@
Refinement was introduced in [LV, Section 5.] . Roughly speaking, refinement stabilises the tight Heegaard splitting by drilling along Legendrian curves embedded in meridional discs. The drilling process cuts each non-product disc into subdiscs, and the curves are chosen so that these subdiscs are product discs for the stabilised Heegaard splitting $\widetilde{\mathcal{H}}$. See Figure 5. This refinement $\widetilde{\mathcal{H}}$ is necessarily convex, so via Lemma 2.5 one may view refinement as a process which associates an open book decomposition to a tight Heegaard splitting of $(M,\xi)$.
Figure 5. To refine a tight Heegaard splitting $\mathcal{H}$, stabilise it via tunnels drilled through non-product discs to produce a convex splitting $\widetilde{H}$. Here, the meriodonal disc $A$ with $|\partial A\cap\Gamma_{\Sigma}|=4$ in $\mathcal{H}$ splits into a pair of product discs $A_{1}$, $A_{2}$ in $\widetilde{\mathcal{H}}$.
@@
[LV, Lemma 3.8 and Theorem 5.10] Suppose that $\widetilde{\mathcal{H}}$ and $\widetilde{\mathcal{H}}^{\prime}$ are refinements of the tight Heegaard splitting $\mathcal{H}$. If $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ are the open book decompositions associated to $\widetilde{\mathcal{H}}$ and $\widetilde{\mathcal{H}}^{\prime}$, respectively, then $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ admit a common positive stabilisation.
@@
Any two convex Heegaard splittings of $(M,\xi)$ admit a common positive stabilisation.
Much of the work of [LV] was devoted to showing that various natural operations one may perform on a tight Heegaard splitting are in fact (sequences of) positive stabilisations.
@@
The proof of Lemma 3.10 in [LV] shows that in a convex Heegaard splitting, attaching any bypass $1$-handle is a positive stabilisation. The following more general statement will be used later.
@@
[LV, Lemma 7.1] Let $\mathcal{H}=(\Sigma,U,V)$ and $\mathcal{H}^{\prime}=(\Sigma^{\prime},U^{\prime},V^{\prime})$ be tight Heegaard splittings of $(M,\xi)$ and assume that $U^{\prime}$ is obtained from $U$ by attaching a single bypass $h^{1}\cup h^{2}$. Let
′′$\mathcal{H}^{\prime\prime}=\big{(}\partial(U\cup h^{1}),U\cup h^{1},M\setminus\text{int}(U\cup h^{1})\big{)}$
be the Heegaard splitting formed by attaching only the $1$-handle associated to this bypass. Then the refinements of $\mathcal{H}$, $\mathcal{H}^{\prime}$ and $\mathcal{H}^{\prime\prime}$ all admit a common positive stabilisations.

### v2

@@ Abstract.
This paper proves the Giroux Correspondence in dimension three using Heegaard splittings of contact manifolds. In [16], the authors proved the Giroux Correspondence for tight contact $3$-manifolds via convex Heegaard surfaces, and simultaneously, [10] gave an all-dimensions proof of the Giroux Correspondence by generalising convex surface theory to higher dimensions. This paper extends the Heegaard splitting approach to arbitrary (not necessarily tight) contact 3-manifolds in order to provide a proof accessible to a low-dimensional audience. The proof assumes a result announced in [10] that classifies moves relating bypass decompositions for isotopic contact structures on cobordisms that are topological products; in the Appendix, we prove this result in the 3-dimensional setting.
@@ 1. Introduction
[Giroux Correspondence [GIR02]] Two open book decompositions support isotopic contact structures if and only if they are related by a sequence of positive (de)stabilisations.
Because topology and contact geometry are so closely entwined, it’s natural to probe whether topological techniques can be squeezed to produce meaningful contact geometric outputs. In this spirit, in 2023 the authors developed new tools to study $3$-dimensional contact manifolds via their Heegaard decompositions. Using this enhanced relationship between contact structures and Heegaard splittings, we proved the Giroux Correspondence for all tight contact $3$-manifolds [16]. Simultaneously, Breen-Honda-Huang used generalised convex surface theory to prove the Giroux Correspondence for all contact structures and in all dimensions [10]. While the work of Breen-Honda-Huang fully covers the $3$-dimensional case, the generality of their techniques lends their proof a different flavour.
This paper provides a complete proof of the Giroux Correspondence in three dimensions, further developing the Heegaard splitting approach to this topic and using an essential result of [10] (Theorem 2.9 below) to bridge the gap to overtwisted manifolds. Not only will this proof be accessible to a low-dimensional audience, but it also illustrates the effectiveness of applying convex surface techniques in concert with Heegaard splittings, an approach which we hope will see more applications in the future. At the request of the referee, we have also included a proof of Theorem 2.9 in Appendix A. Although similar in spirit to the argument in [10], some techniques specific to low dimensions affect the flavour of the argument.
@@ 1.1. Heegaard splittings and open book decompositions
Any open book decomposition of a manifold gives rise to a canonical Heegaard decomposition. Heegaard splittings produced thus are called convex Heegaard splittings, and one may directly recover an open book supporting $\xi$ from any convex Heegaard splitting of $(M,\xi)$. The innovation in [16] is a procedure called refinement that produces an open book from a more general class of Heegaard splittings of a contact manifold. Although this open book decomposition is not uniquely defined, any two open books produced thus are related by a sequence of positive (de)stabilisations – exactly the equivalence appearing in the Giroux Correspondence. It follows that the Giroux Correspondence can be rephrased as a statement about convex Heegaard splittings (Theorem 3.3). Our proof of the Giroux Correspondence for tight contact manifolds relied on showing that many operations one might naturally perform on a Heegaard splitting of a contact manifold in fact preserve the positive stabilisation class of the associated open book.
As explained next, these techniques were insufficient in the case of overtwisted manifolds. In order to assign an open book decomposition to a contact manifold with a Heegaard splitting, we required both handlebodies in the Heegaard splitting to be tight. When two convex Heegaard surfaces are isotopic, there is a sequence of convex Heegaard surfaces interpolating between them. Our approach tried to assign a sequence of open book decompositions to this sequence of Heegaard splittings. However, even when both original Heegaard splittings have tight handlebodies, intermediate Heegaard splittings in this sequence may nevertheless have overtwisted handlebodies. Since we could not preclude this for arbitrary contact $3$-manifolds, our proof of the Giroux Correspondence was limited to tight manifolds.
The present work bypasses this limitation by considering highly stabilised Heegaard splittings where the handlebodies are always tight. We introduce the term “bridging” for a particular method of constructing such a splitting, and we show that different choices made in constructing a bridge splitting preserve the positive stabilisation class of the associated open book. Using bridge splittings, we prove that any pair of convex Heegaard splittings for a fixed contact manifold correspond to open books that are related by a sequence of positive (de)stabilisations.
@@ 1.2. A note on notation
[…]
@@ Acknowledgment
This research was supported in part by the Austrian Science Fund (FWF) P 34318. MS was supported by the Swedish Research Council, VR 2022-06593, Centre of Excellence in Geometry and Physics at Uppsala University and VR 2024-04417, project grant. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. The authors would also like to thank Joseph Breen, Nataliya Goncharuk and Ko Honda for helpful discussions on the topic.
@@ 2. Background
We assume the reader is familiar with basic notions in contact geometry, but we use this section to collect some tools and terminology that will be essential for the rest of the paper. Section 2.2 provides a quick review of convex surface theory and introduces weak contactomorphism (Definition 2.3) as an equivalence relation on contact manifolds with convex boundary. Section 2.3 describes the relationship between open book decompositions and convex Heegaard splittings (Definition 2.6) of a contact manifold. Section 2.4 presents the contact handle model for bypass attachment. Finally, Section 2.5 states the key theorem which classifies distinct bypass decompositions of a product cobordism. This statement will be proved in the A. Also, Lemma 2.10 gives a new perspective on the familiar operation of bypass rotation.
2.1. Legendrian Graphs
In Section 3 will consider contact Heegaard decompositions constructed as neighbourhoods of Legendrian graphs. Working with these Legendrian graphs requires a 1-parameter version of the Legendrian approximation theorem that was originally proved for Legendrian knots in the standard contact structure by Fuchs and Tabachnikov [FT97].
Theorem 2.1.
For $t\in[0,1]$, let $G_{t}$ be a 1-parameter family of spacial graphs in a contact manifold $(M,\xi)$. Then $G_{t}$ can be $C^{0}$-approximated by a 1-parameter family of Legendrian graphs.
That is, for any smooth choice of neighbourhoods $N_{t}$ of $G_{t}$, we can choose a 1-parameter family of Legendrian graphs $L_{t}$ such that $L_{t}$ is smoothly isotopic to $G_{t}$ inside $N_{t}$.
Note that at a vertex $v$ of a Legendrian graph, all the edges must be tangent to the contact plane $\xi_{v}$. It follows that the edges have a cyclic order that must be preserved throughout a Legendrian isotopy. This explains the final hypothesis in the next result.
Corollary 2.2.
Suppose the Legendrian graphs $L_{0}$ and $L_{1}$ are smoothly isotopic, and suppose, too, that the cyclic orders of the edges agree at every vertex. Then after finitely many positive and negative stabilistions, $L_{0}$ and $L_{1}$ become Legendrian isotopic.
Proof of Theorem 2.1.
At each vertex of $G_{t}$, fix a cylic order on the incoming edges. Isotope $G_{t}$ inside $N_{t}$ so that near each vertex $v$, the edges are Legendrian in a Darboux-ball $D_{v}(t)$ and realise the given cyclic order for all $t\in[0,1]$. Now extend the isotopy of $G_{t}$ to an ambient isotopy $\Phi_{t}$ of $M$ that brings the Daurboux balls $D_{v}(0)$ to $D_{v}(t)$ via a contactomorphism. Pull the contact structure back by the ambient isotopy. This produces a fixed graph $G=\Phi^{-1}_{t}(G_{t})$ in a 1-parameter family of contact structures $\xi_{t}=(\Phi_{t})_{*}\xi$, all of which agree near the vertices of $G$. Choose a neighbourhood $N$ of $G$ contained in the intersection of the neighbourhoods $\Phi_{t}^{-1}(N_{t})$.
Now restrict to an edge $e=(u,v)$ and follow the usual procedure for Legendrian approximation, as follows. Using standard transversality arguments, a $C^{\infty}$-small isotopy of $\xi_{t}$ ensures that away from $\coprod D_{v}$, the edge $e$ is transverse to $\xi_{t}$ for all but finitely many times $0<t_{1}<\dots<t_{n}<1$. Furthermore, at each such $t_{j}$, exactly at one point in the interior of $e$ has a non-degenerate tangency with $\xi_{t_{j}}$.
In time intervals away from the singular $t_{j}$ values, take standard neighbourhoods contactomorphic to $(I\times D^{2},\ker(ds+r^{2}d\varphi))$, allowing the radius to vary in $t$. By the compactness of each interval $t_{j}+\varepsilon,t_{j+1}-\varepsilon]$, there exists a radius $R$ defining a neighbouhood contained in $N$ for all $t$. Use a leaf of the characteristic foliation on $T_{R}=I\times\{r=R\}$ together with a Legendrian model in $D_{u}\cup D_{v}$ to approximate $e$ in $N$. Different choices of $R$ correspond to positive (de)stabilisations of the Legendrian approximation.
At a singular times $t_{j}$, the interior of the edge $e$ is tangent to $\xi_{t_{j}}$ at a point $p$. That is, the edges $\Phi_{t-\varepsilon}(e)$ and $\Phi_{t+\varepsilon}(e)$ are related to each other by transverse stabilisation or destabilisation, depending on the sign of $\frac{\mathit{d}}{dt}\alpha_{t}(T_{p}e)|_{t=t_{j}}$. For simplicity, assume $\Phi_{t+\varepsilon}(e)$ is a transverse stabilistion of $\Phi_{t-\varepsilon}(e)$. Then a standard neighbourhood of $\Phi_{t+\varepsilon}(e)$ is contained in the standard neighbourhood of $\Phi_{t-\varepsilon}(e)$, and their Legendrian approximations as above are related to each other via a single negative Legendrian stabilisation and some positive Legendrian (de)stabilistions that depend on the $R$-values chosen for $t_{j-1}+\varepsilon,t_{j}-\varepsilon]$ and $t_{j}+\varepsilon,t_{j+1}-\varepsilon]$.
Thus we have proved that Legendrian isotopy and positive and negative Legendrian stabilisations suffice to approximate $G_{t}$. Corollary 2.2 follows from performing the stabilisations before the isotopies. ∎
@@ 2.2. Convex surfaces
The reader looking for a more thorough introduction to convex surface theory is referred to [MAS14] or the lecture notes [ETN04, 13].
@@
[16, Definition 2.3.] Two contact manifolds $(M,\xi)$ and $(M^{\prime},\xi^{\prime})$ with convex boundary are weakly contactomorphic if there is a contact embedding $\iota\colon M\hookrightarrow M^{\prime}$ such that $\iota(\partial M)$ is convexly isotopic to $\partial M^{\prime}$.
@@
We note that if $\iota:M\rightarrow M^{\prime}$ is a contact embeddeding as above, then $M\setminus\iota({M^{\prime}})$ is a weak identity morphism in the sense of [22].
Henceforth, we will be interested in contact manifolds with boundary only up to weak contactomorphism. Two (weak contactomorphism classes of) manifolds may be glued along diffeomorphic boundary components if and only if there is a diffeomorphism respecting the dividing sets. Gluing is well defined as an operation on weak contactomorphism classes of manifolds; see [16, Proposition 2.8.].
@@
[16, Definition 2.4.] Two embedded codimension-0 submanifolds $N_{1},N_{2}\subset(M,\xi)$ with convex boundary are weakly (contact) isotopic if there is an isotopy $N_{s}$ between them such that $\partial N_{s}$ remains convex throughout.
@@ A Heegaard splitting $(\Sigma,U,V)$ of $(M,\xi)$ is convex if
[16, Proposition 3.4.] Let $(M,\xi)$ be a contact manifold and let $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ be open books supporting $\xi$. Then the induced convex Heegaard splittings are isotopic via an isotopy keeping the Heegaard surface convex if and only if $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ are isotopic through a path of open books supporting $\xi$.
@@ 2.4. Bypass attachment
Figure 1. Left: a bypass half-disc attached along the arc $c$. Centre: the attached 11-handle is a regular neighbourhood of an arc Legendrian isotopic to $c$. Right: attaching the shaded 22-handle along $c\cup c^{\prime}$ completes the bypass attachment.
The steps described here are illustrated in Figure 1. First, attach a minimally twisting contact $1$-handle $h^{1}$ along the 0-sphere $\partial c$. Slightly abusing notation, continue to denote the remaining portion of the bypass arc by $c$. Let $c^{\prime}\subset\partial h^{1}$ be a Legendrian arc with endpoints on $c$ and such that ${\mathit{tb}}(c\cup c^{\prime})=-1$. Attach a contact $2$-handle $h^{2}$ along $c\cup c^{\prime}$. Topologically, the handles cancel, but in general the dividing set on the new boundary is not isotopic to $\Gamma_{\Sigma}$. See Figure 2. We often write $X$ to denote the pair of handles $h^{1}\cup h^{2}$; $X$ is itself a neighbourhood of a bypass half-disc $D$ whose cornered Legendrian boundary consists of the bypass arc $c$ and a Legendrian core of $h^{1}$. Coordinate models for this construction are provided in [OZB11], and the dividing sets at each stage of the construction are shown in Figure 1.
@@ 2.5. Bypass decompositions
$X_{1}\ast\dots\ast X_{j}\ast X_{j+1}\ast\dots\ast X_{f}$$X_{1}\ast\dots\ast X_{j+1}\ast X_{j}\ast\dots\ast X_{f}$
The main theorem of both [22] and [10] establishes that Trival Insertion and Far Commutation suffice to relate any two bypass decompositions of topologically trivial products:
@@
[10, Theorem 3.1.2 ] Any two bypass decompositions of $(\Sigma\times I,\xi)$ are related by Bypass Arc Isotopy and a finite iteration of Trivial Insertion and Far Commutation moves.
The dimension 3 case was the Main Theorem in [22], although presented in different language. For completeness, we give a proof of this statement using independent techniques in the Appendix.
@@
Figure 4. Attach a $1$-handle $h_{T}^{1}$ along a push-off of the Legendrian curve $d$. Then isotope the disc cobounded by $h_{T}^{1}$ and $d$ across the product disc AA to produce a bypass disc $D$ cobounded by $h_{T}^{1}$ and $c_{T}$.
@@ 3. Heegaard Splittings and the Giroux Correspondence
With useful tools and vocabulary established, we now turn to proving the Giroux Correspondence. As a first step, Section 3.1 summarises results from [16] that allow us to restate the theorem about a pair of open books as a theorem about a pair of Heegaard splittings. Section 3.2 introduces bridging as an operation to stabilise a Heegaard splitting of a contact manifold. After establishing some properties of bridge splittings, we relate them to refinements in Section 3.3 to prove the main result.
@@ 3.1. Tight Heegaard splittings and refinement
As described in Section 2.3, each convex Heegaard splitting of a contact manifold corresponds to an open book decomposition supporting the contact structure. Here we recall a key idea from [16] that extends the class of Heegaard splittings of a contact manifold which determine open books.
@@
Refinement was introduced in [16, Section 5.] . Roughly speaking, refinement stabilises the tight Heegaard splitting by drilling along Legendrian curves embedded in meridional discs. The drilling process cuts each non-product disc into subdiscs, and the curves are chosen so that these subdiscs are product discs for the stabilised Heegaard splitting $\widetilde{\mathcal{H}}$. See Figure 5. This refinement $\widetilde{\mathcal{H}}$ is necessarily convex, so via Lemma 2.7 one may view refinement as a process which associates an open book decomposition to a tight Heegaard splitting of $(M,\xi)$.
Figure 5. To refine a tight Heegaard splitting $\mathcal{H}$, stabilise it via tunnels drilled through non-product discs to produce a convex splitting $\widetilde{H}$. Here, the meriodonal disc $A$ in $\mathcal{H}$ with $|\partial A\cap\Gamma_{\Sigma}|=4$ splits into a pair of product discs $A_{1}$, $A_{2}$ in $\widetilde{\mathcal{H}}$.
@@
[16, Lemma 3.8 and Theorem 5.10] Suppose that $\widetilde{\mathcal{H}}$ and $\widetilde{\mathcal{H}}^{\prime}$ are refinements of the tight Heegaard splitting $\mathcal{H}$. If $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ are the open book decompositions associated to $\widetilde{\mathcal{H}}$ and $\widetilde{\mathcal{H}}^{\prime}$, respectively, then $(B,\pi)$ and $(B^{\prime},\pi^{\prime})$ are related by a sequence of positive (de)stabilisations. .
@@
Any two convex Heegaard splittings of $(M,\xi)$ are related by a sequence of positive (de)stabilisations.
Much of the work of [16] was devoted to showing that various natural operations one may perform on a tight Heegaard splitting are in fact (sequences of) positive stabilisations.
@@
The proof of Lemma 3.10 in [16] shows that in a convex Heegaard splitting, attaching any bypass $1$-handle is a positive stabilisation. The following more general statement will be used later.
@@
[16, Lemma 7.1] Let $\mathcal{H}=(\Sigma,U,V)$ and $\mathcal{H}^{\prime}=(\Sigma^{\prime},U^{\prime},V^{\prime})$ be tight Heegaard splittings of $(M,\xi)$ and assume that $U^{\prime}$ is obtained from $U$ by attaching a single bypass $h^{1}\cup h^{2}$. Let
$\mathcal{H}^{\prime\prime}=\big(\partial(U\cup h^{1}),U\cup h^{1},M\setminus\text{int}(U\cup h^{1})\big)$
be the Heegaard splitting formed by attaching only the $1$-handle associated to this bypass. Then the refinements of $\mathcal{H}$, $\mathcal{H}^{\prime}$ and $\mathcal{H}^{\prime\prime}$ are all related by sequences of positive (de)stabilisations.